# On stable assignments generated by choice functions of mixed type

### v1

@@ Abstract
We consider one variant of stable assignment problems in a bipartite graph endowed with nonnegative capacities on the edges and quotas on the vertices. It can be viewed as a generalization of the stable allocation problem introduced by Bai̋ou and Balinsky, which arises when strong linear orders of preferences on the vertices in the latter are replaced by weak ones. At the same time, our stability problems can be stated in the framework of a theory by Alkan and Gale on stable schedule matchings generated by choice functions of a wide class. In our case, the choice functions are of a special, so-called mixed, type. The main content of this paper is devoted to a study of rotations in our mixed model, functions determining “elementary” transformations of one stable assignments into others; they look more sophisticated compared with rotations in the stable allocation problem that are generated by simple cycles. On this way, we explicitly construct a weighted poset of rotations and show that the stable assignments are bijective to the so-called closed functions for this poset.
@@ 1 Introduction
Since the famous work of Gale and Shapley [7] on stable matchings (“marriages”) in bipartite graphs, there have appeared a variety of researches on models of stability on graphs. A well-known model among them was introduced by Bai̋ou and Balinsky [2] under the name of stable allocation problem.
@@
In a setting of that problem, there is given a bipartite graph $G=(V,E)$ whose vertices are partitioned into two subsets (color classes), say, $F$ and $W$, interpreted as the sets of firms and workers. An edge $e\in E$ connecting vertices (or “agents”) $u$ and $v$ represents a possible “contract” between them, and its intensity (or an assignment on $e$) is restricted by a nonnegative real upper bound (or capacity) $b(e)\in{\mathbb{R}}_{+}$. Besides, each vertex $v\in V$ is endowed with a quota $q(v)\in{\mathbb{R}}_{+}$ which restricts the total assignment on the set $E_{v}$ of edges incident to $v$, and there is a (strong) linear order $>_{v}$ establishing preferences on the edges in $E_{v}$, namely, $e>_{v}e^{\prime}$ means that “agent” $v$ prefers “contract” $e$ to “contract” $e^{\prime}$. An assignment (or allocation) $x:E\to{\mathbb{R}}_{+}$ satisfying the capacity and quota constraints is called stable if, roughly, there is no edge $e=\{u,v\}$ such that $x(e)$ can be increased, without violation of the capacities and quotas anywhere, possibly decreasing $x$ on an edge $e^{\prime}\in E_{u}$ with $e^{\prime}<_{u}e$ and/or an edge $e^{\prime\prime}\in E_{v}$ with $e^{\prime\prime}<_{v}e$.
@@
Bai̋ou and Balinsky [2] proved that a stable allocation always exists and can be found by an efficient algorithm. Also they showed that if the capacities and quotas are integer-valued, then there exists an integral stable allocation, and obtained some other impressive results. (Hereinafter an algorithm dealing with $G=(V,E)$ is said to be efficient if it runs in time polynomial in $|E|$, assuming w.l.o.g. that $|E|\geq|V|-1$.)
@@
Alkan and Gale [1] considered much more general problem in a bipartite graph with capacities $b$ and, possibly, quotas $q$ as above. Here for each $v\in V$, there is a choice function $C_{v}$ that maps the set ${\cal B}_{v}:=\{z\in{\mathbb{R}}_{+}^{E_{v}}\colon z(e)\leq b(e)\;\forall e\in E_{v}\}$ into itself. The role of $C_{v}$ is twofold; in the case with quotas, this can be briefly explained as follows. First, when the amount $|z|:=\sum(z(e)\colon e\in E_{v})$ exceeds the quota $q(v)$, it reduces $z$ so as to get the equality $|z^{\prime}|=q(v)$, where $z^{\prime}=C_{v}(z)$. Second, it gives a way to compare the vectors in ${\cal B}_{v}$ obeying the quotas, by use of the preference relation $\succeq\,=\,\succeq_{v}$ defined by $z\succeq z^{\prime}\Longleftrightarrow C_{v}(z\vee z^{\prime})=z$ (where $z\vee z^{\prime}$ takes the values $\max\{z(e),z^{\prime}(e)\}$). By imposing to the choice functions $C_{v}$, $v\in V$, the axioms of consistence and persistence (which have been known, sometimes under different names, for simpler models, going back to Kelso and Crawford [13], Roth [14], Blair [3], and some others) and adding to them the axiom of size-monotonicity (cf. Fleiner [6]) or, stronger, of quota-filling, they succeeded to build a rather powerful and elegant theory of stability for weighted bipartite graphs. A central result in it (generalizing earlier results for particular cases) is that the set ${\cal S}$ of stable assignments (or stable schedule matchings, in terminology of [1]) is nonempty and represents a distributive lattice under the preference relations $\succeq_{F}$ determined by the choice functions $C_{v}$ with $v$ contained in the vertex part $F$ of $G$; namely, $x,y\in{\cal S}$ satisfy $x\succeq_{F}y$ if and only if for each $f\in F$, the restrictions $z,z^{\prime}$ of $x,y$ to $E_{v}$ (respectively) satisfy $z\succeq_{f}z^{\prime}$.
@@
As an illustration in the setting with quotas, Alkan and Gale pointed out two special cases (Examples 1,2 in Sect. 2 of [1]). One of them is just the above-mentioned stable allocation problem, that we will briefly denote as SAP. In this case, for each $v\in V$, the choice function $C_{v}$ acts on $z\in{\cal B}_{v}$ with $|z|>q(v)$ so as to respect more preferred and ignore less preferred edges in $E_{v}$; more precisely, it selects the element $\widehat{e}\in E_{v}$ such that $p:=\sum(z(e)\colon e>_{v}\widehat{e})<q(v)$ and $p+z(\widehat{e})\geq q(v)$, preserves the values for all $e>_{v}\widehat{e}$, reduces to zero the values for all $e<_{v}\widehat{e}$, and changes the value on $\widehat{e}$ to $q(v)-p$. The second special case deals with the so-called diversifying choice function for each $v\in V$, which uniformly cuts values of $z\in{\cal B}_{v}$ with $|z|>q(v)$ from above; more precisely, it computes the number (“level”) $r$ such that $\sum(\min\{z(e),r\}\colon e\in E_{v})=q(v)$, and changes $z(e)$ to $\min\{z(e),r\}$, $e\in E_{v}$. In fact, this corresponds to the equivalence (or disorder) relation between all elements of $E_{v}$ from the viewpoint of the “agent” $v$, in contrast to the allocation case where $E_{v}$ is totally ordered. Surprisingly, the diversifying model is poor: for any $b$ and $q$, there exists exactly one stable solution, see [12].
@@
In this paper, we consider a common generalization of SAP and the stable diversifying problem. It is obtained from SAP by admitting that for each $v\in V$, the preferences on $E_{v}$ are established by a weak linear order. Equivalently, the set $E_{v}$ is partitioned into subsets $\pi^{1},\ldots,\pi^{k}$, called ties, such that $e\in\pi^{i}$ prefers $e^{\prime}\in\pi^{j}$ if $i<j$, and $e,e^{\prime}$ are regarded as equivalent if $i=j$. We write $e>_{v}e^{\prime}$ if $i<j$, and $e\sim_{v}e^{\prime}$ if $i=j$. Then an assignment $x:E\to{\mathbb{R}}_{+}$ satisfying the capacities and quotas is stable if, roughly, there is no edge $e=\{u,v\}$ such that $x(e)$ can be increased, without violation of the capacities and quotas anywhere, possibly decreasing $x$ on an edge $e^{\prime}\in E_{u}$ such that either $e^{\prime}<_{u}e$, or $e^{\prime}\sim_{u}e$ and $x^{\prime}(e)>x(e)$, and/or an edge $e^{\prime\prime}\in E_{v}$ such that either $e^{\prime\prime}<_{v}e$, or $e^{\prime\prime}\sim_{v}e$ and $x^{\prime\prime}(e)>x(e)$. This case is within the framework of Alkan-Gale’s model as well; here for $v\in V$, the choice function $C_{v}$ respects the assignments on more preferred ties, ignore less preferred ties, and acts as in the diversifying case for one tie in between those (a precise definition will be given later). We refer to a $C_{v}$ of this sort as a choice function of mixed type, and denote the arising stable assignment problem as SMP.
We will take advantages from both ways to state SMP: via weak linear orders (a combinatorial setting), and via choice functions of mixed type (an analytic setting). On the second way, relying on Alkan-Gale’s theory, we utilize the fact that the set ${\cal S}={\cal S}^{\rm mix}(G,b,q,\geq)$ of stable assignments in SMP is nonempty and forms a distributive lattice under the corresponding preference relations on ${\cal S}$, appeal to the construction of the $F$-optimal stable assignment, that we denote as $x^{\rm min}$, and use some other results from [1].
The main content of this paper is devoted to a study of rotations, “elementary augmentations” connecting close stable assignments in SMP, which clarifies, to some extent, the structure of the lattice on ${\cal S}$. This is already beyond the content of [1], but is parallel, in a sense, to a study of rotations for SAP given by Dean and Munshi [4]. (Note that originally nice properties of rotations and their applications have been revealed in connection with the stable marriage and roommate problems in [8, 9, 10]. One can also mention a recent work [5] on the Boolean variant of Alkan-Gale’s model, where rotations and their posets are extensively studied.)
It turns out that the formation and properties of rotations in SMP are more intricate compared with SAP. In particular, while a rotation in SAP is generated by a simple cycle whose edges have unit weights, a rotation in SMP is represented by an integer-valued function $\rho$ on the edges of a certain strongly connected subgraph $L$, and it is possible that $\rho$ takes large values and $L$ has large vertex degrees. As a consequence, while SAP with integral $b,q$ always has an integral stable solution, SMP with integral $b,q$ does not guarantee the existence of a stable solution with small denominators; moreover, the convex hull of ${\cal S}$ may have vertices with exponentially large denominators. Nevertheless, we can explicitly construct, in time polynomial in $|E|$, the set of all possible rotations $\rho$ in SMP, along with their maximal possible weights $\tau(\rho)$ (that can be used in “elementary” transformations $x\rightsquigarrow x^{\prime}=x+\tau(\rho)\rho$ of stable assignments). Moreover, by an analogy with SAP, we can efficiently construct a poset of rotations $\rho$ endowed with weights $\tau(\rho)$ and show that the so-called closed functions associated with this poset one-to-one correspond to the stable assignments in SMP (thus obtaining a sort of “explicit representation” for ${\cal S}$).
This paper is organized as follows. Section 2 contains basic definitions and settings and reviews important results from [1] needed to us. Here we recall that in a general case: the set ${\cal S}$ of stable assignments is nonempty; the partial orders $\succeq_{F}$ and $\succeq_{W}$ on ${\cal S}$ that express preferences for the parts $F$ (“firms”) and $W$ (“workers”) are polar to each other, i.e., for $x,y\in{\cal S}$, if $x\succeq_{F}y$ then $y\succeq_{W}x$, and vice versa; and $({\cal S},\succeq_{F})$ is a distributive lattice. In this lattice, we denote the best assignment for $F$ (and the worst for $W$) by $x^{\rm min}$, and the worst assignment for $F$ by $x^{\rm max}$. In Section 3 we recall the general method from [1] to construct a sequence of assignments that converges to $x^{\rm min}$. This method may take infinitely many iterations even if we deal with our model SMP. In the Appendix we devise an efficient modification to find $x^{\rm min}$ in SMP.
Section 4 is devoted to an in-depth study of rotations and related aspects in SMP. Here for the lattice $({\cal S},\succeq_{F})$ and an element $x\in{\cal S}-\{x^{\rm max}\}$, we consider the set ${\cal C}(x)$ of vectors (“elementary augmentations”) $\Delta\in{\mathbb{R}}^{E}$ such that $|\Delta|$ is sufficiently small and $x^{\prime}:=x+\Delta$ is a stable assignment satisfying $x\succ_{F}x^{\prime}$. The set ${\cal C}(x)$ is of conic type, in the sense that $\Delta\in{\cal C}(x)$ and $0\leq\varepsilon\leq 1$ imply $\varepsilon\Delta\in{\cal C}(x)$, and the irreducible vectors in ${\cal C}(x)$ can be characterized combinatorially. More precisely, we associate with $x$ a certain directed graph $\overrightarrow{\Gamma}(x)$ whose underlying undirected graph is a subgraph of $G$, which possesses the following property: each maximal strongly connected component of $\overrightarrow{\Gamma}(x)$ is bijective to an irreducible vector in ${\cal C}(x)$ considered up to proportionality. (Here a strongly connected, or, simply, strong component $L$ of $\overrightarrow{\Gamma}(x)$ is called maximal if no other strong component of $\overrightarrow{\Gamma}(x)$ is reachable by a directed path from $L$.) We denote by ${\cal L}^{\rm rot}(x)$ the set of maximal strong components in $\overrightarrow{\Gamma}(x)$, and associate with each $L=(V_{L},E_{L})\in{\cal L}^{\rm rot}$ a certain integer vector $\rho_{L}\in{\mathbb{Z}}^{E}$ having the support $E_{L}$ and representing a sort of integer circulation.
This vector $\rho_{L}$ is just what we mean under the rotation for $x$ generated by $L\in{\cal L}^{\rm rot}(x)$. One shows that the vectors $\varepsilon\rho_{L}$, $L\in{\cal L}^{\rm rot}(x)$, with a sufficiently small $\varepsilon>0$, are exactly the irreducible vectors in ${\cal C}(x)$ (cf. Corollary 4.7). (Note that in SAP each rotation $\rho$ is generated by a simple cycle and $\rho$ takes values $0,\pm 1$. In contrast, in SMP maximal strong components $L$ in $\overrightarrow{\Gamma}(x)$ may be rather cumbersome and $\rho_{L}$ may take big values; an example is illustrated in Sect. 5.) For $L\in{\cal L}^{\rm rot}(x)$, we define $\tau(\rho_{L})$ to be the maximal value $\lambda\geq 0$ such that $x+\lambda\rho_{L}$ is stable, and prove that for any $0<\lambda\leq\tau(\rho_{L})$, the assignment $x^{\prime}:=x+\lambda\rho_{L}$ is stable as well (cf. Lemma 4.8). We say that such an $x^{\prime}$ is obtained from $x$ by the rotational shift of weight $\lambda$ along $\rho_{L}$; when $\lambda=\tau(\rho_{L})$, the shift is called full. Another important fact shown in this section (generalizing a known analog for SAP) is that $x$ can be updated by using several rotations simultaneously; namely, for any $\lambda:{\cal L}^{\rm rot}(x)\to{\mathbb{R}}_{+}$ satisfying $\lambda(L)\leq\tau_{L}$ for all $L\in{\cal L}^{\rm rot}(x)$, the assignment $x^{\prime}:=x+\sum(\lambda(L)\rho_{L}\colon L\in{\cal L}^{\rm rot}(x))$ is stable, and $x\succeq_{F}x^{\prime}$ (cf. Proposition 4.9).
Section 5 discusses additional, including computational, aspects of rotations. We explain that for $x\in{\cal S}-\{x^{\rm max}\}$, the set of rotations $\rho=\rho_{L}$ applicable to $x$ (where $L\in{\cal L}^{\rm rot}$) and their maximal weights $\tau(\rho)$ can be found in time polynomial in $|E|$ (note that to construct the function $\rho_{L}$ from the component $L$, we have to solve a system of linear equalities). Also one shows that any $x\in{\cal S}$ can be reached from $x^{\rm min}$ by a sequence of $O(|E|)$ rotational shifts. The section finishes with an example (mentioned above) of rotation $\rho_{L}$ taking “big” values while $L$ is “small”.
@@
In the Appendix, we elaborate an efficient algorithm to find the $F$-optimal stable assignment $x^{\rm min}$ in SMP. It modifies the method from Sect. 3, based on aggregation of “long” sequences of iterations in the latter, by solving certain linear programs.

### v2

@@ Abstract
We consider one variant of stable assignment problems in a bipartite graph endowed with nonnegative capacities on the edges and quotas on the vertices. It can be viewed as a generalization of the stable allocation problem introduced by Bai̋ou and Balinsky, which arises when strong linear orders of preferences on the vertices in the latter are replaced by weak ones. At the same time, our stability problem can be stated in the framework of a theory by Alkan and Gale on stable schedule matchings generated by choice functions of a wide scope. In our case, the choice functions are of a special, so-called mixed, type.
The main content of this paper is devoted to a study of rotations in our mixed model, functions on the edges determining “elementary” transformations between close stable assignments. These look more sophisticated compared with rotations in the stable allocation problem (which are generated by simple cycles). We efficiently construct a poset of rotations and show that the stable assignments are in bijection with the so-called closed functions for this poset; this gives rise to a “compact” affine representation for the lattice of stable assignments and leads to an efficient method to find a stable assignment of minimum cost.
@@ 1 Introduction
Since the famous work of Gale and Shapley [9] on stable matchings (“marriages”) in bipartite graphs, there have appeared a variety of researches on stability models in graphs. A well-known model among them was introduced by Bai̋ou and Balinsky [2] under the name of stable allocation problem.
@@
In a setting of that problem, there is given a bipartite graph $G=(V,E)$ whose vertices are partitioned into two subsets (color classes), say, $F$ and $W$, interpreted as the sets of firms and workers. An edge $e\in E$ connecting vertices (or “agents”) $u$ and $v$ represents a possible “contract” between them, and its intensity (or an assignment on $e$) is restricted by a nonnegative real upper bound, or capacity, $b(e)\in{\mathbb{R}}_{+}$. Besides, each vertex $v\in V$ is endowed with a quota $q(v)\in{\mathbb{R}}_{+}$ which restricts the total assignment on the set $E_{v}$ of edges incident to $v$, and there is a (strong) linear order $>_{v}$ establishing preferences on the edges in $E_{v}$, namely, $e>_{v}e^{\prime}$ means that “agent” $v$ prefers “contract” $e$ to “contract” $e^{\prime}$. An assignment (or allocation) $x:E\to{\mathbb{R}}_{+}$ satisfying the capacity and quota constraints is called stable if, roughly, there is no edge $e=\{u,v\}$ such that $x(e)$ can be increased, without violation of the capacities and quotas anywhere, possibly simultaneously decreasing $x$ on an edge $e^{\prime}\in E_{u}$ with $e^{\prime}<_{u}e$ and/or an edge $e^{\prime\prime}\in E_{v}$ with $e^{\prime\prime}<_{v}e$.
@@
Bai̋ou and Balinsky [2] proved that a stable allocation always exists and can be found by an efficient algorithm. Also they showed that if the capacities and quotas are integer-valued, then there exists an integral stable allocation, and obtained some other impressive results. (Hereinafter an algorithm dealing with $G=(V,E)$ is said to be efficient if it runs in time polynomial in $|V|,|E|$.)
@@
Alkan and Gale [1] considered much more general problem in a bipartite graph with capacities $b$ and, possibly, quotas $q$ as above. Here each vertex $v\in V$ is endowed with a choice function $C_{v}$ that maps the set ${\cal B}_{v}:=\{z\in{\mathbb{R}}_{+}^{E_{v}}\colon z(e)\leq b(e)\;\forall e\in E_{v}\}$ into itself. The role of $C_{v}$ is twofold; in the case with quotas, this can be briefly explained as follows. First, when the amount $|z|:=\sum(z(e)\colon e\in E_{v})$ exceeds the quota $q(v)$, it reduces $z$ so as to get the equality $|z^{\prime}|=q(v)$, where $z^{\prime}=C_{v}(z)$. Second, it gives a way to compare the vectors in ${\cal B}_{v}$ obeying the quotas, by use of the preference relation $\succeq\,=\,\succeq_{v}$ defined by $z\succeq z^{\prime}\Longleftrightarrow C_{v}(z\vee z^{\prime})=z$ (where $z\vee z^{\prime}$ takes the values $\max\{z(e),z^{\prime}(e)\}$). By imposing to the choice functions $C_{v}$, $v\in V$, the axioms of consistence and persistence (or substitutability) (which have been known, sometimes under different names, for simpler models, going back to Kelso and Crawford [16], Roth [20], Blair [5], and some other works) and by adding to them the axiom of size-monotonicity (cf. Fleiner [8]) or, stronger, of quota-filling, Alkan and Gale developed a rather powerful and elegant theory of stability for weighted bipartite graphs. A central result in it (generalizing earlier results for particular cases) is that the set ${\cal S}$ of stable assignments (or stable schedule matchings, in terminology of [1]) is nonempty and forms a distributive lattice under the preference relations $\succeq_{F}$ determined by the choice functions $C_{v}$ with $v$ contained in the vertex part $F$ of $G$. (Here $x,y\in{\cal S}$ satisfy $x\succeq_{F}y$ if and only if for each $f\in F$, the restrictions $z,z^{\prime}$ of $x,y$ to $E_{v}$ (respectively) satisfy $z\succeq_{f}z^{\prime}$.)
@@
As an illustration of the setting with quotas, Alkan and Gale pointed out two special cases (Examples 1,2 in Sect. 2 of [1]). One of them is just the above-mentioned stable allocation problem, that we will briefly denote as SAP. In this case, for each $v\in V$, the choice function $C_{v}$ acts on $z\in{\cal B}_{v}$ with $|z|>q(v)$ so as to respect more preferred and ignore less preferred edges in $E_{v}$. More precisely, $C_{v}$ selects the element $\widehat{e}\in E_{v}$ such that $p:=\sum(z(e)\colon e>_{v}\widehat{e})<q(v)$ and $p+z(\widehat{e})\geq q(v)$, preserves the values for all $e>_{v}\widehat{e}$, reduces to zero the values for all $e<_{v}\widehat{e}$, and changes the value on $\widehat{e}$ to $q(v)-p$. The second special case, that we abbreviate as SDP, deals with the so-called diversifying choice function for each $v\in V$, which uniformly reduces values of $z\in{\cal B}_{v}$ with $|z|>q(v)$; more precisely, it computes the number (“height”) $r$ such that $\sum(\min\{z(e),r\}\colon e\in E_{v})=q(v)$, and changes $z(e)$ to $\min\{z(e),r\}$, $e\in E_{v}$. In fact, this corresponds to the equivalence (or disorder) relation between all elements of $E_{v}$ from the viewpoint of the “agent” $v$, in contrast to the allocation case where $E_{v}$ is totally ordered. A rather surprising fact shown in [15] is that SDP is poor: for any $b$ and $q$, there exists exactly one stable solution.
@@
In this paper, we consider a common generalization of SAP and SDP, which is interesting theoretically and may find reasonable applications. It differs from SAP by admitting that for each $v\in V$, the preferences on $E_{v}$ are established by a weak linear order. Equivalently, the set $E_{v}$ is partitioned into subsets $\pi^{1},\ldots,\pi^{k}$, called ties, such that $v$ prefers $e\in\pi^{i}$ to $e^{\prime}\in\pi^{j}$ if $i<j$, and regards $e,e^{\prime}$ as equivalent if $i=j$. We write $e>_{v}e^{\prime}$ if $i<j$, and $e\sim_{v}e^{\prime}$ if $i=j$. Then an assignment $x:E\to{\mathbb{R}}_{+}$ satisfying the capacity and quota constraints as before is stable if, roughly, there is no edge $e=\{u,v\}$ such that $x(e)$ can be increased, without violation of the capacities and quotas anywhere, possibly simultaneously decreasing $x$ on an edge $e^{\prime}\in E_{u}$ such that either $e^{\prime}<_{u}e$ and $x(e^{\prime})>0$, or $e^{\prime}\sim_{u}e$ and $x^{\prime}(e)>x(e)$, and/or an edge $e^{\prime\prime}\in E_{v}$ such that either $e^{\prime\prime}<_{v}e$ and $x(e^{\prime\prime})>0$, or $e^{\prime\prime}\sim_{v}e$ and $x^{\prime\prime}(e)>x(e)$. Our case is shown to be within the framework of Alkan-Gale’s model as well; here for $v\in V$, the choice function $C_{v}$ respects the assignments on more preferred ties, ignore less preferred ties, and acts as in the diversifying case for one tie in between those (a precise definition will be given later). We refer to a $C_{v}$ of this sort as a choice function of mixed type, and denote the arising stable assignment problem as SMP.
@@
We will take advantages from both ways to state SMP: via weak linear orders (a combinatorial setting), and via choice functions of mixed type (an analytic setting). On the second way, relying on Alkan-Gale’s theory in [1], we utilize the fact that the set ${\cal S}={\cal S}^{\rm mix}(G,b,q,\geq)$ of stable assignments in SMP is nonempty and forms a distributive lattice under the corresponding preference relations on ${\cal S}$, and appeal to the construction of the $F$-optimal stable assignment, that we denote as $x^{\rm min}$.
The main content of this paper is devoted to a study of rotations, “elementary augmentations” connecting close stable assignments in SMP, which leads to clarifying the structure of the lattice on ${\cal S}$. This is already beyond the content of [1], but is parallel, in a sense, to a study of rotations for SAP in [6]. (Note that originally nice properties of rotations and their applications have been revealed in connection with the stable marriage and roommate problems in [11, 12, 13]. One can also mention a recent work [7] on the Boolean case of Alkan-Gale’s model (generalizing stable matchings of many-to-many type), where rotations are extensively studied.)
It turns out that the formation and properties of rotations in SMP are more intricate compared with SAP. Whereas a rotation in SAP is generated by a simple cycle whose edges have weights ±$\pm 1$, a rotation in SMP is represented by an integer-valued function $\rho$ on the edges of a certain strongly connected graph $L$, and it is possible that $\rho$ takes large values and $L$ has large vertex degrees. As a consequence, SMP with integral $b,q$ need not have an integral stable solution (in contrast to SAP); moreover, the convex hull of ${\cal S}$ may have vertices with exponentially large denominators.
Nevertheless, we can explicitly construct, in time polynomial in $|E|$, the set of all possible rotations $\rho$ in SMP, along with their maximal possible weights $\tau(\rho)$ (that can be used in “elementary” transformations $x\rightsquigarrow x+\tau(\rho)\rho$ of stable assignments). Moreover, we efficiently construct a poset of rotations $\rho$ endowed with weights $\tau(\rho)$ and show that the so-called closed functions related to this poset one-to-one correspond to the stable assignments in SMP, yielding a sort of “explicit and compact representation” for ${\cal S}$.
We then take advantages from this bijection between the set ${\cal S}$ of stable assignments in SMP and the set $\mathfrak{C}$ of closed functions for the rotational poset; this is in parallel with the corresponding results on the above-mentioned stability problems in [6, 7] and earlier results on rotations and their applications in stable matching problems of one-to-one and many-to-many types in [12, 13, 3]. Using the fact that the number of rotations is $O(|E|)$, one can solve, in strongly polynomial time, the weighted strengthening of stable assignment problem: given weights (or costs) $c(e)\in{\mathbb{Z}}$ of edges $e\in E$, find $x\in{\cal S}$ minimizing $cx:=\sum(c(e)x(e)\colon e\in E)$ (similar to results in [13, 3, 6]). This is due to a reduction to the problem of minimizing a linear function over $\mathfrak{C}$, which in turn is reduced to the classical minimum cut problem, by Picard’s method [19].
As an additional observation, note that once we are able to minimize a linear function over ${\cal S}$ in polynomial time, the separation problem over ${\cal S}$ is polynomially solvable as well, due to the polynomial equivalence between linear optimization and separation problems established in [10] based on the ellipsoid method.
(Recall that the separation problem for ${\cal S}$ is: given $\xi\in{\mathbb{R}}^{E}$, decide whether $\xi$ belongs to the convex hull ${\rm conv}({\cal S})$ of ${\cal S}$, and if not, find a hyperplane in ${\mathbb{R}}^{E}$ separating $\xi$ from ${\cal S}$. Note that since the separation problem for ${\cal S}$ is efficiently solvable, it is tempting to ask of a possibility to explicitly characterize facets of the convex hull ${\rm conv}({\cal S})$ of ${\cal S}$. A characterization of this sort is well known for the set of stable marriages in a bipartite graph, due to Vande Vate [23]. Another example of availability of a good facet characterization concerns the Boolean case of Alkan–Gale’s model for which [7] shows that the convex hull of stable objects is affinely congruent to an appropriate Stanley’s order polytope. One more example is the unconstrained case $b\equiv\infty$ of SAP for which an affine congruence to an order polytope is shown in [18]. As to our SMP model, the facet characterization issue is open even in the unconstrained case.)
Concerning computational complexity aspects, it should be noted that all algorithms that we present in this paper are, in fact, strongly polynomial, i.e. (roughly) they use standard arithmetic and logic operations, and the number of these is bounded by a polynomial in the amount (not the size) of input data; in a sense, they are adjusted to work with real $b,q$ as well. Moreover, when dealing with integer-valued $b,q$, the memory size needed for storing intermediate data should be bounded by a polynomial in the size of input data. Note that we often do not care of precisely estimating time bounds of the algorithms that we devise and restrict ourselves merely by establishing their (strongly) polynomial-time complexity; equivalently, we may say that an algorithm takes time polynomial in $|V|,|E|$.
This paper is organized as follows. Section 2 contains basic definitions and settings and reviews important results from [1] needed to us. Here we recall that in a general case: the set ${\cal S}$ of stable assignments is nonempty; the partial orders $\succeq_{F}$ and $\succeq_{W}$ on ${\cal S}$ that express preferences for the parts $F$ (“firms”) and $W$ (“workers”) are polar to each other, i.e., for $x,y\in{\cal S}$, if $x\succeq_{F}y$ then $y\succeq_{W}x$, and vice versa; and $({\cal S},\succeq_{F})$ is a distributive lattice. In this lattice, we denote the best assignment for $F$ (and the worst for $W$) by $x^{\rm min}$, and the worst assignment for $F$ by $x^{\rm max}$. In Section 3 we recall the general method from [1] to construct a sequence of assignments that converges to $x^{\rm min}$. This method may take infinitely many iterations even for our SMP model. For this reason, in the Appendix we devise an efficient modification to find $x^{\rm min}$ in SMP.
Section 4 is devoted to an in-depth study of rotations and related aspects in SMP. Here for the lattice $({\cal S},\succeq_{F})$ and an element $x\in{\cal S}-\{x^{\rm max}\}$, we consider the set $X^{\varepsilon}(x)$ of stable assignments $x^{\prime}$ satisfying $x\succ_{F}x^{\prime}$ and $|x-x^{\prime}|:=\sum(|x(e)-x^{\prime}(e)|\colon e\in E)=\varepsilon$ for a sufficiently small real $\varepsilon>0$. The set $X^{\varepsilon}(x)$ forms a polytope whose vertices can be characterized combinatorially. More precisely, we construct a certain directed graph $\overrightarrow{\Gamma}(x)$ such that the maximal strong (viz. strongly connected) components in it are in bijection with the vertices of $X^{\varepsilon}(x)$. (Here a strong component $L$ of $\overrightarrow{\Gamma}(x)$ is called maximal if no other strong component of $\overrightarrow{\Gamma}(x)$ is reachable by a directed path from $L$.) We denote by ${\cal L}^{\rm rot}(x)$ the set of maximal strong components in $\overrightarrow{\Gamma}(x)$ and associate with each $L=(V_{L},E_{L})\in{\cal L}^{\rm rot}$ a certain integer vector $\rho_{L}\in{\mathbb{Z}}^{E}$ having the support $E_{L}$ and forming an integer circulation.
This vector $\rho_{L}$ is just what we mean under the rotation for $x$ generated by $L\in{\cal L}^{\rm rot}(x)$. One shows that the vectors $x+\lambda\rho_{L}$, where $L\in{\cal L}^{\rm rot}(x)$ and $\lambda=\varepsilon/|\rho_{L}|$, are exactly the vertices of $X^{\varepsilon}(x)$ (cf. Corollary 4.9). For each $L\in{\cal L}^{\rm rot}(x)$, we define $\tau(\rho_{L})$ to be the maximal value $\lambda\in{\mathbb{R}}_{+}$ such that $x+\lambda\rho_{L}$ is stable. For $0<\lambda\leq\tau(\rho_{L})$, we say that the assignment $x^{\prime}:=x+\lambda\rho_{L}$ (which is shown to be stable) is obtained from $x$ by the rotational shift of weight $\lambda$ along $\rho_{L}$; when $\lambda=\tau(\rho_{L})$, the shift is called full.
Section 5 discusses additional, including computational, aspects of rotations. We explain that for $x\in{\cal S}-\{x^{\rm max}\}$, the set of rotations $\rho=\rho_{L}$ applicable to $x$ (i.e. $L\in{\cal L}^{\rm rot}$) and their maximal weights $\tau(\rho)$ can be found in time polynomial in $|E|$ (by solving a certain linear system). Also one shows that any $x\in{\cal S}$ can be reached from $x^{\rm min}$ by a sequence of $O(|E|)$ rotational shifts. The section finishes with an example of rotation $\rho_{L}$ taking “big” values while $L$ is “small”; this is extended to a series of rotations whose norms grow exponentially in the number of vertices (in Remark 3).
[…]
Based on these results, in Section 7, we discuss an explicit affine representation $x=x^{\rm min}+A\lambda$ (where $A$ is an $E\times{\cal R}$ matrix) involving stable assignments $x$ and closed functions $\lambda$, and finish with the minimum cost strengthening of SMP. Also we briefly discuss (in Remark 5) interrelations between optimization, separation and facet description issues concerning SMP and earlier stability problems.
In the Appendix, we describe an efficient algorithm to find the $F$-optimal stable assignment $x^{\rm min}$ in SMP. It modifies the method from Sect. 3, by aggregating “long” sequences of iterations in the latter; this is based on solving a certain linear program. Also, as a partial alternative, we present (in Remark 6) an efficient and relatively simple “combinatorial” algorithm which, given an instance of SMP, recognizes the situation when a stable assignment must attain quotas with equalities in all vertices, and if so, finds an optimal stable assignment. The Appendix finishes with an open question.